# ThermoHands: A Benchmark for 3D Hand Pose Estimation from Egocentric Thermal Images

### v4

@@ Abstract.
* Equal contribution
† Corresponding author: xiaoxuan.lu@ucl.ac.uk
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06copyright: none
@@ 1. Introduction
Building on the above insights, this study probes the following research question: Can egocentric thermal imagery be effectively used for 3D hand pose estimation under various conditions (such as different lighting and handwear), and how does it compare to techniques using RGB, NIR, and depth [^1] For readability, we treat depth and NIR images as two ‘spectra’, despite their usual overlap. spectral imagery? To answer this, we introduce ThermoHands  [^2] Anonymous project page: https://thermohands.github.io/, the first benchmark specifically tailored for egocentric 3D hand pose estimation utilizing thermal imaging. This benchmark is supported by a novel multi-spectral and multi-view dataset designed for egocentric 3D hand pose estimation and is unique in comprising thermal, NIR, depth, and RGB images (cfFig. 1). Our dataset emulates real-world application contexts by incorporating both hand-object and hand-virtual interaction activities, with participation from 28 subjects to ensure a broad representation of actions. To offer a thorough comparison across spectral types, we gather data under five distinct scenarios, each characterized by varying environments, handwear, and lighting conditions (cfTab. 1). Considering the challenges associated with manually annotating large-scale 3D hand poses, we developed an automated annotation pipeline. This pipeline leverages multi-view RGB and depth imagery to accurately and efficiently generate 3D hand pose ground truths through optimization based on the MANO model (MANO, ) (cfFig. 4).
@@
    We introduce the first-of-its-kind benchmark, dubbed ThermoHands, to investigate the potential of thermal imaging for egocentric 3D hand pose estimation.
@@
    We will release our dataset, code and models and maintain the benchmark to serve as a new challenge in 3D hand pose estimation.
@@ 3.1. Multi-Spectral Hand Pose Dataset
Overview. At the core of our benchmark lies a multi-spectral dataset for 3D hand pose estimation (cfTab. 1), capturing hand actions performed by 28 subjects of various ethnicities and genders [^3] The study has received the ethical approval from XXX (no name for blind review), and participant consent forms were signed before the collection.. As shown in Fig. 1, we develop a customized head-mounted sensor platform (HMSP) and an exocentric platform to record multi-view data. During the capture, our participants are asked to perform pre-defined hand-object and hand-virtual interaction actions within the playground above the table. The main part is captured in the normal office scenario. To facilitate the evaluation under different settings, four auxiliary parts are recorded i) under the darkness, ii) under the sun glare, iii) with gloves on hand, and iv) in the kitchen environment with different actions, respectively.
@@ 5. TherFormer: A Baseline Method
For our benchmark, we setup a baseline method, dubbed TherFormer, for thermal image-based 3D hand pose estimation. As exhibited in Fig. 5, TherFormer features in its two consecuative transformer modules to model the spatio-temporal relationship of hand joints while being computationally efficient. Note that TherFormer is also capable of processing other spectral images due to their format consistency with thermal images. However, our primary objective is to establish a baseline method for future research, rather than focusing exclusively on methods for thermal images.
@@
     (a)      A2J-Transformer (jiang2023a2j, )      Single      51.68      0.474      20.76      0.603      34
     (b)      HTT (wen2023hierarchical, )      Single      49.09      0.489      20.69      0.599      211
     (c)      TherFormer-S      Single      44.64      0.539      18.34      0.643      136
     (d)      (c) w/o spatial transformer      Single      48.79      0.491      20.15      0.609      174
     (e)      (c) w/o mask guidance      Single      48.83      0.494      18.89      0.625      141
     (f)      HTT (wen2023hierarchical, )      Sequence      47.07      0.512      17.49      0.659      129
     (g)      TherFormer-V      Sequence      43.36      0.549      17.36      0.661      52
Table 3. Comparison between TherFormer and state of the arts on thermal camera-based 3D hand pose estimation. The fps indicates the number of inference steps that models can run per second.
@@ 6.1. Evaluation of the Annotation Method
As the first step, we validate the accuracy of our 3D hand pose annotation (cfSec. 4) and analyze the impact for optimization results. For evaluation, we manually annotate two random sequences from our main dataset, with a total of over 600 frames. To that end, we first annotate the 2D joint locations on RGB images from two viewpoints and obtain the 3D joint positions by triangulation. We calculate the average 3D joint errors across all frames to measure the accuracy.
@@ 6.2. Experiment Setup
Method and Implementation. To provide sufficient baselines for follow-up works, we selected two state-of-the-art methods in 3D hand pose estimation: HTT (wen2023hierarchical, ) and A2J-Transformer (jiang2023a2j, ), and reproduced them on our dataset for experiments. HTT (wen2023hierarchical, ) is a video-based method thus enabling the evaluation in both two problem settings while A2J-Transformer (jiang2023a2j, ) only works under the single image-based setting. For a feasible comparison, we use the same sequence length, i.e., $T=8$, for HTT and TherFormer-V baselines. We exclude the additional action block used by HTT (wen2023hierarchical, ), focusing solely on hand pose estimation. We adjusted the anchor initialization phase of A2J-Transformer (jiang2023a2j, ) to better accommodate our dataset, without altering the density of its anchors. All trained models are tested on a single NVIDIA RTX 4090 GPU to fairly compare their inference speed.
Evaluation Metrics. Following HTT (wen2023hierarchical, ), we evaluate the accuracy of 3D hand pose estimation with two metrics: Percentage of Correct Keypoints (PCK) and Mean End-Point Error (MEPE) (zimmermann2017learning, ), in both camera space and root-aligned (RA) space. For the RA space, we align the estimated wrist with its groundtruth position before measuring the joint errors. In terms of PCK, we report the corresponding Area Under the Curve (AUC) over the 0-50mm/80mm error thresholds for the camera/RA space.
@@ 6.3. Thermal Image-based Results
Main Results. To assess TherFormer’s performance in thermal image-based 3D hand pose estimation, we compare it against state-of-the-art methods (wen2023hierarchical, ; jiang2023a2j, ) on the main testing set, as shown in Tab. 3. TherFormer-S outforms two competing methods under the single image-based setting, while TherFormer-V surpasses the counterpart HTT (wen2023hierarchical, ) given the same sequential images as input. Such an improvement mainly stems from our mask-guided spatial attention design that can adaptively encode the spatial interaction among hand joints with the guidance of the hand mask (cfSec. 5). A performance gap can be observed between single image-based and video-based settings for both HTT (wen2023hierarchical, ) and TherFormer. We credit this to their usage of temporal information that helps to tackle the occlusion cases and solve ambiguities. Thanks to our lightweight network design, TherFormer is highly efficient to run with fps of 136 and 52 for single and sequence input respectively, ensuring its real-time application to resource-constrained devices. Moreover, TherFormer-V also improves the performance over HTT (wen2023hierarchical, ) for other spectra (cfTab. 5), proving its adaptability to different inputs.
@@ 7. Conclusion
Limitation. As the first exploration in this field, this work has certain limitations that can guide our future work. For example, we only annotate 3D hand pose for our dataset, which limits its usage of evaluation to other tasks relevant to human hands. Further efforts could be annotating the fine-grained hand action splits and hand-object contact (brahmbhatt2020contactpose, ).
@@
We also provide more visualization in the format of both images and videos for our data, annotation and results on our anonymous project page.

### v5

@@ Abstract.
3D Hand Pose Estimation, Thermal Vision, Hand Pose Dataset
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: The 23rd ACM Conference on Embedded Networked Sensor Systems; May 06–09, 2025; Irvine, USAbooktitle: The 23rd ACM Conference on Embedded Networked Sensor Systems (SenSys’25), May 06–09, 2025, Irvine, USAprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing → Human computer interaction (HCI)ccs: Computer systems organization → Embedded and cyber-physical systemsccs: Human-centered computing → Ubiquitous and mobile computing
@@ 1. Introduction
Building on the above insights, this study probes the following research question: Can egocentric thermal imagery be effectively used for 3D hand pose estimation under various conditions (such as different lighting and handwear), and how does it compare to techniques using RGB, NIR, and depth [^1] For readability, we treat depth and NIR as two ‘spectra’, despite their usual overlap. spectral imagery? To answer this, we introduce ThermoHands  [^2] Project page: https://thermohands.github.io/., the first benchmark specifically tailored for egocentric 3D hand pose estimation utilizing thermal imaging. This benchmark is supported by a novel multi-spectral and multi-view dataset designed for egocentric 3D hand pose estimation and is unique in comprising thermal, NIR, depth, and RGB images (cfFig. 1). Our dataset emulates real-world application contexts by incorporating both hand-object and hand-virtual interaction activities, with participation from 28 subjects to ensure a broad representation of actions. To offer a thorough comparison across spectral types, we gather data under five distinct scenarios, each characterized by varying environments, handwear, and lighting conditions (cfTab. 1). Considering the challenges associated with manually annotating large-scale 3D hand poses, we developed an automated annotation pipeline. This pipeline leverages multi-view RGB and depth imagery to accurately and efficiently generate 3D hand pose ground truths through optimization based on the MANO model (MANO, ) (cfFig. 4).
@@
    We introduce the first-of-its-kind benchmark, ThermoHands, to investigate the potential of thermal imaging for egocentric 3D hand pose estimation.
@@
    We release our dataset, code and models and maintain the benchmark at https://github.com/LawrenceZ22/ThermoHands.
@@ 3.1. Multi-Spectral Hand Pose Dataset
Overview. At the core of our benchmark lies a multi-spectral dataset for 3D hand pose estimation (cfTab. 1), capturing hand actions performed by 28 subjects of various ethnicities and genders [^3] The study has received the ethical approval from University of Edinburgh, and participant consent forms were signed before the collection.. As shown in Fig. 1, we develop a customized head-mounted sensor platform (HMSP) and an exocentric platform to record multi-view data. During the capture, our participants are asked to perform pre-defined hand-object and hand-virtual interaction actions within the playground above the table. The main part is captured in the normal office scenario. To facilitate the evaluation under different settings, four auxiliary parts are recorded i) under the darkness, ii) under the sun glare, iii) with gloves on hand, and iv) in the kitchen environment with different actions, respectively.
@@ 5. TherFormer: A Baseline Method
For our benchmark, we setup a baseline method, dubbed TherFormer, for thermal image-based 3D hand pose estimation. As exhibited in Fig. 5, TherFormer features in its two consecutive transformer modules to model the spatio-temporal relationship of hand joints while being computationally efficient. Note that TherFormer is also capable of processing other spectral images due to their format consistency with thermal images. However, our primary objective is to establish a baseline method for future research, rather than focusing exclusively on methods for thermal images.
@@
     (a)      HaMeR* (pavlakos2024reconstructing, )      Single      -      -      20.88      0.598      118
     (b)      A2J-Transformer (jiang2023a2j, )      Single      51.68      0.474      20.76      0.603      34
     (c)      HTT (wen2023hierarchical, )      Single      49.09      0.489      20.69      0.599      211
     (d)      TherFormer-S      Single      44.64      0.539      18.34      0.643      136
     (e)      (c) w/o spatial transformer      Single      48.79      0.491      20.15      0.609      174
     (f)      (c) w/o mask guidance      Single      48.83      0.494      18.89      0.625      141
     (g)      HTT (wen2023hierarchical, )      Sequence      47.07      0.512      17.49      0.659      129
     (h)      TherFormer-V      Sequence      43.36      0.549      17.36      0.661      52
Table 3. Comparison between TherFormer and state of the arts on thermal camera-based 3D hand pose estimation. The fps indicates the number of inference steps that models can run per second. HaMeR* estimates 3D hand pose in a root-aligned manner by default; thus, we can only evaluate MEPE-RA and AUC-RA metrics.
@@ 6.1. Evaluation of the Annotation Method
As the first step, we validate the accuracy of our 3D hand pose annotation (cfSec. 4) and analyze the impact on optimization results. For evaluation, we manually annotate two random sequences from our main dataset, with a total of over 600 frames. To that end, we first annotate the 2D joint locations on RGB images from two viewpoints and obtain the 3D joint positions by triangulation. We calculate the average 3D joint errors across all frames to measure the accuracy.
@@ 6.2. Experiment Setup
Method and Implementation. To provide sufficient baselines for follow-up works, we selected three state-of-the-art methods in 3D hand pose estimation: HTT (wen2023hierarchical, ), A2J-Transformer (jiang2023a2j, ) and HaMeR (pavlakos2024reconstructing, ), and reproduced them on our dataset for experiments. HTT (wen2023hierarchical, ) is a video-based method thus enabling the evaluation in both two problem settings while A2J-Transformer (jiang2023a2j, ) only works under the single image-based setting. HaMeR (pavlakos2024reconstructing, ) is a model-based method that takes a single image as input and reconstructs a MANO-based hand mesh. It estimates each hand individually from a cropped image patch centered on the hand, removing the need for explicit localization. For comparison, we extract 3D keypoints from its predicted mesh as the estimated pose. We use the same sequence length, i.e., $T=8$, for HTT and TherFormer-V baselines. We exclude the additional action block used by HTT (wen2023hierarchical, ), focusing solely on hand pose estimation. We adjusted the anchor initialization phase of A2J-Transformer (jiang2023a2j, ) to better accommodate our dataset, without altering the density of its anchors. All trained models are tested on a single NVIDIA RTX 4090 GPU to fairly compare their inference speed.
Evaluation Metrics. Following HTT (wen2023hierarchical, ), we evaluate the accuracy of 3D hand pose estimation with two metrics: Percentage of Correct Keypoints (PCK) and Mean End-Point Error (MEPE) (zimmermann2017learning, ), in both camera space and root-aligned (RA) space. For the RA space, we align the estimated wrist with its groundtruth position before measuring the joint errors. In terms of PCK, we report the corresponding Area Under the Curve (AUC) over the 0-80mm/50mm error thresholds for the camera/RA space.
@@ 6.3. Thermal Image-based Results
Main Results. To assess TherFormer’s performance in thermal image-based 3D hand pose estimation, we compare it against state-of-the-art methods (wen2023hierarchical, ; jiang2023a2j, ) on the main testing set, as shown in Tab. 3. TherFormer-S outforms three competing methods under the single image-based setting, while TherFormer-V surpasses the counterpart HTT (wen2023hierarchical, ) given the same sequential images as input. Such an improvement mainly stems from our mask-guided spatial attention design that can adaptively encode the spatial interaction among hand joints with the guidance of the hand mask (cfSec. 5). A performance gap can be observed between single image-based and video-based settings for both HTT (wen2023hierarchical, ) and TherFormer. We credit this to their usage of temporal information that helps to tackle the occlusion cases and solve ambiguities. Thanks to our lightweight network design, TherFormer is highly efficient to run with fps of 136 and 52 for single and sequence input respectively, ensuring its real-time application to resource-constrained devices. Moreover, TherFormer-V also improves the performance over HTT (wen2023hierarchical, ) for other spectra (cfTab. 5), proving its adaptability to different inputs.
@@ 6.4.3. Quantitative Results under Normal Conditions
7. Limitation and Future Work
As the first exploration in this field, this work has certain limitations that can guide our future work. First, our dataset primarily focuses on controlled indoor environment and challenging scenarios for RGB/NIR spectrum. Data collection in other challenging scenarios, especially where thermal imaging can be beneficial or limited, is currently scarce. In future works, we plan to expand our dataset beyond the current settings to incorporate a wider range of real-world challenges, including grasping hots objects, reflective surfaces to thermal radiation, ambient heating objects, and more outdoor environments. Paired data should be collected by including and removing these challenging factors to quantify their impact to thermal imaging-based 3D hand pose estimation. Second, we only annotate 3D hand pose for our dataset, which limits its usage of evaluation to other tasks relevant to human hands. Further efforts could be annotating the fine-grained hand action splits and hand-object contact (brahmbhatt2020contactpose, ).
@@ 8. Conclusion
Acknowledgments
This research is partially supported by the Engineering and Physical Sciences Research Council (EPSRC) under the Centre for Doctoral Training in Robotics and Autonomous Systems at the Edinburgh Centre of Robotics (EP/S023208/1) and the grants from the Cisco Research.
@@ Acknowledgments
We also provide more visualization in the format of both images and videos for our data, annotation and results on our project page.
@@ Appendix B Dataset Annotation
Usage Limitations. Thanks to the well-established RGB-based tools like (kirillov2023segment, ; mediapipe_hands_2020, ), we can extract accurate 2D hand information, which, combined with depth images, enable MANO-based 3D hand pose annotation. However, this approach relies on adequate illumination and minimal occlusion, as hands may appear degraded in RGB or depth images under challenging conditions such as gloves, darkness, and sun glare, and RGB-based tools lack robustness against these factors. Moreover, while our annotation pipeline provides accurate 3D hand pose labels, it is impractical for real-time egocentric estimation due to its reliance on multi-view input, high computational cost, and iterative MANO optimization. In contrast, 3D hand pose estimation methods are designed for real-time inference from a single image view, more suitable for practical deployment.